When 0.055 mol of propionic acid, C2H5CO2H, is dissolved in 550 mL of water, the equilibrium concentration of H3O+ ions is measured to be 1.71 x 10−3 M. What is Ka for this acid?
Answer #1

[C2H5CO2H] = mol of C2H5CO2H / volume in L

= 0.055 mol / 0.550 L

= 0.1 M

Given at equilibrium:

[H+] = 1.71*10^-3 = x

Use:

Ka = [C2H5CO2-][H+] / [C2H5CO2H]

= x*x / (0.1625-x)

= (1.71*10^-3)*(1.71*10^-3) / (0.1 - 1.71*10^-3)

= 2.97*10^-5

Answer: 2.97*10^-5